. If astronauts could travel at v = 0.950c, we on Earth would say it takes (4.20/0.950) = 4.42 years to reach Alpha Centauri, 4.20 light-years away. The astronauts disagree.

    (a) How much time passes on the astronauts’ clocks?

    (b) What is the distance to Alpha Centauri as measured by the astronauts?


1
Expert's answer
2022-01-25T06:50:19-0500

(a) Time will be dilated by a factor of γ in astronaut's frame.

Time

"t = t_0 \times γ \\

 γ = \sqrt{1 - (\frac{v}{c})^2 } \\

= \sqrt{1 - (\frac{0.950c }{ c})^2 } \\

= \sqrt{0.0975} \\

= 0.312"

time as measured by astronaut "= 4.42 \times 0.312 = 1.379 \; years"

(b) Distance measured "= L_0 \times γ = 4.2 \times 0.312 = 1.31 \; light \;years"